A 30kg girl is riding a 20kg bike with a velocity of 10/m/s due east. What is the total momentum of a girl and a bike together?
1
Expert's answer
2019-03-14T17:34:35-0400

By the definition of the momentum we have:


"p_{tot} = (m_{girl} + m_{bike})v = (30 kg + 20 kg) \cdot 10 \dfrac{m}{s} = 500 kg \cdot \dfrac{m}{s}."

Answer:

"p_{tot} = 500 kg \cdot \dfrac{m}{s}."